evening shift summary, 7/1/99

AG (ag@fnal.gov)
Mon, 5 Jul 1999 00:36:35 -0500 (CDT)

Shift Summary for eve shift, 7/4/99

On Shift: BobT, AG

Hours of beam available at all intensities: 3.2
Hours of beam available at nominal intensity: 2.5
Hours of E832 data (special runs not included): 3.2
Hours of special runs (laser scans, muon runs...): 0.

Accounting of lost time:

Tevatron's down.

========================================================================

Runs ended this shift: none ( 13816 is still running ).

Tapes brought to FCC this shift: none.

========================================================================

Accesses during shift:

1. 16:40, bad MU3V PS removed.
2. 18:00, ETOT LV PS checked and trimmed.

========================================================================

Overview of shift:

Mostly no beam, but normal E832 datataking during the beam time.
2 short accesses were made during the down time.

========================================================================

Description of Shift:

run____spill_____time______comment

------------------------------------------------------------------------

13816__150_______1540______Stream 5 missing ffff word -> kfrend

13816__157_______1550______ktev3 hangs, Ed is working on it from home ..
___________________________Turns out the problem was with DDD veto on
___________________________plane 3. Re-loading current maps for busy
___________________________logic fixes it.

13816__174_______1613______Laser stab rms for one channel is 10%.
___________________________Goes away during the next spill.

_______178_______1617______Tev injection problems... Investigating.
___________________________Est. 20 min.
___________________________We are getting a spill every 5-6 minutes...
___________________________We decide to make an access to replace dead
___________________________MU3X PS, hoping that we can use H PS from
___________________________VV'H as a spare... Tough luck, V' and H use
___________________________the same PS. We simply removed the bad PS,
___________________________(note that this FLUKE supply was at -2250).
___________________________MCR promises beam shortly.

_______225_______1725______Instead of beam we get the message:
___________________________Attempting to regain MI PS regulation;
___________________________problems are heat related. Experts enroute.
___________________________No estimate at this time.

_______243_______1740______Called MCR, the estimate is 30-45 min.

_______248_______1750______Sydney passes by. At the same time Etotal
___________________________button on SDAQ monitoring window goes red:
___________________________3 thresholds are higher than normal.
___________________________Sydney fixes it.

_______254_______1800______Sydney and Arun go on 20 min access.
___________________________They checked and trimmed a little LV on
___________________________the ETOT PS which was swapped earlier this
___________________________week.

_______304_______1900______No beam. The message is still the same...

_______350_______2000______Still no beam. Called MCR, they don't have
___________________________an estimate yet. The experts think they'll
___________________________be able to tell more during the next hour.

_______406_______2123______MCR calls and promises beam....

_______407_______2125______And here it comes!

_______420_______2143______One laser stab rms is 10%. It's channel 757.
___________________________Goes away during the next spill.

_______517_______2348______Unknown crate in stream 5 -> kfrend.

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

Page Elizabeth for BA base repair only if beam down for
at least 2 hours.